### v2

# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observe that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 4.9 times more efficiently on average compared to baseline agents.111[Uncaptioned image] Code: pull-ups/VisEscape
1 Introduction
Realistic goals often necessitate exploration; for pantry cooking, one must search through cabinets to identify available ingredients and determine how they can be combined into a meal. Likewise, a trip to a foreign city cannot be planned without searching for major sights and train tables. In such cases, effective planning depends on acquiring information through interaction with the environment rather than having it provided a priori.
Existing works on embodied agents have addressed the importance of exploration, but it has primarily been studied under decomposable and explicit tasks like “Place a cleaned egg in a microwave” [38, 32, 21, 18]. However, the capability of agents to autonomously search for and utilize relevant information from environments where tasks or solutions are not explicitly defined has not been sufficiently investigated.
Figure 1: In VisEscape, player actions cause state changes that may or may not be immediately visible. For instance, pressing a button might move an object located outside the player’s immediate field of view. These key characteristics—partial observability and dynamic state changes—require players to maintain spatial awareness and continually update their memory.
Figure 2: An illustration of an excerpt of a trajectory from one room in VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes in different locations to solve creative puzzles.
To bridge this gap, we argue that practical evaluation of decision-making agents should incorporate exploration as a core component. To elicit meaningful exploration, we claim that two key criteria must be met: (1) the state should be partially observable at each timestep, meaning that the agent can only perceive limited aspects of the environment based on its current position and sensors, rather than having full knowledge of the surroundings; and (2) the setup should necessitate a dynamic update of the agent’s state knowledge in response to its actions, ensuring that the effects of those actions—such as modifications to the environment or newly available actions resulting from the acquisition of a new object or removal of constraints—might not be directly observable at the current viewpoint, requiring the agent to actively investigate and continually adapt.
To evaluate agents under these conditions, we propose VisEscape, a benchmark of 20 virtual escape rooms for assessing exploration-centric problem-solving capacities of multimodal agents. Given the implicit and unspecified goal of “escape the room”, the agents explore their surroundings, identify and integrate relevant information, discover how objects can be used, and solve complex puzzles to successfully escape. Moreover, they need to retain previously observed information and interpret indirect clues, requiring them to reason beyond immediate observations and dynamically adapt to environmental changes.
Additionally, VisEscape provides observation as an image for each state and presents executable actions in text form. This brings together the advantages of text-based games, which can define diverse action types (e.g., pouring coffee from a pot or cutting a chain with a nipper), and the benefit of being able to focus on evaluating the agent’s planning or reasoning capabilities of diverse VLMs by isolating the execution of actions from low-level action control.
On VisEscape, we evaluate various VLMs’ performance and observe that even advanced models like Claude-3.5-Sonnet and GPT-4o show low escape rates of only 10%, while all other models achieve 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Feedback, and ReAct [45]—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escaping successfully escaped, while models with pre-existing escape capabilities demonstrate $\times$3.7 enhanced effectiveness and $\times$4.9 improved efficiency. Additionally, we discover that frequent failures of low-performing agents involve repetition of meaningless actions, and that their deficiencies can be mitigated with enhanced reasoning capacity.
In summary, the key points of this paper are threefold:
  • We introduce VisEscape, a unique and challenging benchmark for evaluating exploration-centric problem-solving abilities of multimodal agents in dynamic environments.
  • We propose VisEscaper framework, which significantly improves multimodal agents’ performance in terms of effectiveness and efficiency.
  • We address the deficiencies in reasoning capacity of VLMs by leveraging the reasoning capacity of LLMs, thereby offering valuable insights into potential improvements for multimodal agents.
2 Related Works
2.1 Interactive Environments for Evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world [8, 17, 33]. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge [40] and ethical decision-making [31].
Several studies have advanced LLM/VLM-based agents built on Minecraft [39, 25, 11, 16]. In parallel, embodied agent simulators have facilitated evaluating language-driven agents interacting within visually grounded environments [19, 32, 37, 27].
Meanwhile, the GUI environment has recently been proposed as a benchmark [10, 9, 41] for evaluating VLMs’ interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfaces[24, 7], reason about their functionality [34], and execute actions to successfully complete given tasks [35, 13, 44, 49].
2.2 LLM/VLM as an Action Decision Maker
Deploying LLM or VLM agents in real-world applications requires effective handling of dynamic and unpredictable environments [29]. Such environments are characterized by changing contexts, unexpected events, and continuous novel observations [30, 47]. To successfully adapt, agents must actively self-evolve using mechanisms for adaptive reasoning and continuous learning [50].
Recent studies have highlighted three critical capabilities for agent adaptation. First, accumulation of experience enables agents to remember previous interactions, particularly failures, and use these experiences to improve future decisions [12]. Second, strategic exploration allows agents to intelligently choose actions based on past outcomes and external feedback, balancing the benefits of learning against potential risks [51]. Third, knowledge abstraction helps agents form generalized high-level representations from raw observations, significantly improving their adaptability in diverse situations [48].
To support these adaptation skills, we integrate three specialized modules into our agent architecture; Memory, Feedback, and ReAct [45]. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
3 VisEscape Construction
3.1 Key Characteristics of VisEscape
1. Self-directed successive task. In all rooms of VisEscape, the only instruction is “escape the room”, without any guidance on how to progress. Under this uncertainty about where to begin, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires an average of 27.2 steps to successfully escape, even for the oracle trajectory. Agents must independently explore, discovering objectives through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement, with earlier discoveries becoming relevant later in the process.
2. Diverse interaction with objects. Since all actions available in VisEscape are defined as text, it was possible to define numerous diverse and creative actions that fit with objects placed in the room. For example, an agent can discover a key hidden inside bread by cutting the bread with a knife, open a closet by cutting chains with nippers, and check a hidden code by flipping an hourglass. These interactions are determined by the type of object currently being viewed and the items currently possessed. When an agent performs an action that satisfies the room’s predefined conditions, it triggers state transitions within the room.
Figure 3: Examples of visual quizzes that requires associative thinking in VisEscape. For all problems, we controlled the spatial positioning of the locks and clue images to ensure they remain distinctly separated. This design forces the agent not only to store each piece of information independently but also to identify relevant connections between them (such as “color” in the example above) and reason based on the established associations between these separate pieces of information.
Figure 4: An illustration of each component in VisEscape.
Scene Receptacle Item Quiz Interaction
# Total 1084 118 102 27 311
# Types 662 32 23 27 37
Avg./room 54.2 5.9 5.1 1.4 15.6
Table 1: Statistics for scenes, receptacles, items, quizzes, and interactions in VisEscape.
3. Associative visual quizzes. One key element of escape rooms is finding connections between different pieces of information, which requires associative thinking—the ability to link seemingly unrelated elements and a fundamental component of creativity [28, 5, 4]. For example, associative thinking helps players connect a symbol on a puzzle with a similar symbol elsewhere and recognize how objects sharing the same color might be related. As escape rooms are ideal for evaluating the associative thinking abilities of VLMs—an area that has remained unexplored—we placed one or two number/letter code locks and corresponding hand-crafted visual quizzes in each room, which require such reasoning. See Figure 3 for examples.
3.2 Dataset Composition
VisEscape consists of 20 distinct rooms. Each room comprises four walls corresponding to four directions, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in [38]. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room contains an average of 5.9 receptacles and 5.1 items, along with one or two visual quizzes. Statistics and examples are shown in Table 1 and Figure 4.
3.3 Game Graph
We construct a game graph to model the state of each room in VisEscape providing visual observations corresponding to the current state of the game. The agent can “inspect [object]” to examine receptacles or items more closely by switching to a close-up view, navigate around the room by using “turn to [direction]”, and interact with objects by performing actions defined by each object. Each room has an average of 54.2 unique scenes, which change depending on the actions performed.
3.4 Room Creation Process
To design the room, we utilize three different 3D modeling software programs: Trimble Sketchup, Autodesk Revit, and Chaos Enscape. Specifically, Trimble Sketchup is utilized for object design, Autodesk Revit for object placement and the overall scene composition, and Chaos Enscape for photorealistic rendering. Refer to the Appendix C. for the detailed dataset construction process.
Figure 5: An overview of the three modules comprising VisEscaper, along with examples of inputs and outputs for each module.
4 VisEscaper: Exploring Agent in VisEscape
Figure 5 provides an overview of the VisEscaper framework, which comprises three modules: the Memory, Feedback, and ReAct [45] modules. Each module, along with task-specific instructions, selectively receives information from the environment and responses from other modules, allowing it to efficiently perform tasks by only considering the information essential for that specific module. For the Memory module, since directly using images from multiple turns of observations as input can exceed the model’s context-length capacity, we utilize the same VLM as an image captioner to replace visual observations with textual descriptions.
4.1 Memory module
Since the agent can only observe a limited portion of the room, which causes the entire state of the room to be partially observable, the agent must be equipped with memory capabilities in order to make progress. VisEscaper’s memory consists of three components:
  • Structured spatial memory: This memory hierarchically stores information about room walls, the receptacles on those walls, the states of these receptacles, along with their visual information and characteristics. This serves as a directional aid for navigating toward specific receptacles or items.
  • Exploration memory: This memory stores information about which objects and items have been inspected so far and which have not. This encourages the agent to explore unobserved information rather than getting stuck on the information acquired thus far.
  • Salient action memory: This memory selectively stores actions that are worth remembering, such as interactions with objects, attempting to enter passwords, etc. This allows the agent to make action decisions by referencing actions it has performed in the past.
The Memory module of VisEscaper serves a crucial function in mitigating the computational constraints imposed by extended logs during trajectories and addressing the inherent context window constraints of the model while organizing verbose information into structured and prioritized data.
4.2 Feedback module
The Feedback module takes the action performed by the agent along with the preceding and subsequent observations, allowing it to independently analyze the outcomes or impacts of its actions. As illustrated in Figure 5, even if the observations remain unchanged before and after the action, the Feedback module can conclude that “$0$ is not the correct password for this lock.”, which would be difficult to obtain from a single observation alone. Such feedback enables the agent to determine its current actions based on the analysis of the results of its past actions.
4.3 ReAct module
The ReAct module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research [45], the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition, the agent transmits its reasoning from the previous step to the next step, enabling it to maintain consistent short-term plans across multiple steps.
5 Experiments
[…]
Claude-3.5-Sonnet VisEscaper 21.7 (x3.44) 32.37 (x2.91) 10.92 (x3.26) 249.5 (x0.87)
BaseAgent 6.3 11.11 3.35 286.7
GPT-4o VisEscaper 13.3 (x4.03) 29.02 (x1.81) 3.91 (x6.51) 276.2 (x0.94)
BaseAgent 3.3 16.00 0.60 294.5
GPT-4o-mini VisEscaper 6.7 21.31 (x3.97) 1.05 291.2 (x0.97)
BaseAgent 0.0 5.37 0.0 300.0
Random 0.0 0.90 0.0 300.0
Table 2: Comparison of random baseline, BaseAgent, and VisEscaper for Claude-3.5, GPT-4o, and GPT-4o-mini. The highest improvement ratios achieved by VisEscaper compared to BaseAgent for each metric are bolded, and the highest scores of the BaseAgent on each metric are underlined.
5.1 Experiment Setup
For each room, we conducted three different trials and averaged the resulting metrics. To prevent a few challenging puzzles from blocking access to subsequent checkpoints, we experimented with both the default no-hint setting ($exp_{\text{base}}$) and the hint-provided setting ($exp_{\text{hint}}$). In $exp_{\text{hint}}$, if there are no changes in achieved checkpoints for over 30 consecutive turns, a hint guideline message is given to guide the agent toward the next checkpoint. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded after a total of 300 steps.
$exp_{\text{base}}$ $exp_{\text{hint}}$
[…]
Claude 3.5 Sonnet 21.7 32.37 10.92 249.5 90.0 96.33 18.87 30.46 103.1
GPT-4o 13.3 29.02 3.91 276.2 95.0 97.39 20.74 26.83 112.9
GPT-4o-mini 6.7 21.31 1.05 291.2 46.7 67.40 32.16 8.93 228.2
InternVL2.5-38B 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2-40B 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
LLaVA-v1.6-34B 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-OneVision-7B 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
Human 82.5 95.80 51.30 52.8
Random 0.0 0.90 0.00 -
Table 3: The results of diverse VLMs applied to VisEscaper across both $exp_{\text{base}}$ and $exp_{\text{hint}}$. For each experimental setting and metric, the best scores are bolded and the second-best scores are underlined. Experiments without step counts all ended in failure due to reaching the maximum turn limit. For human scores, we assigned four rooms per participant, with a total of 20 participants in the $exp_{\text{base}}$ setting. Refer to Appendix D. for detailed information on the human study.
5.2 Metrics
We adopted the major metrics widely used in benchmarks for embodied agents.
  • Success Rate (SR): The ratio of escape success in an episode
  • Goal Completion (GC): The ratio of checkpoints completed at the end of an episode
  • Step: Total actions taken by an agent.
  • Success weighted by Path Length [2] (SPL): Success rate weighted by below term:
    Path length of the oracle trajectoryPath length taken by the $\frac{\text{Path length of the oracle trajectory}}{\text{Path length taken by the agent}}$
    This term penalizes the success rate based on the total number of steps taken.
For $exp_{\text{hint}}$, we define a metric regarding the agent’s hint usage: Hint-assisted Completion Rate (HCR), which denotes the ratio of checkpoints completed by the agent with the assistance of hints. Higher is better for all metrics except for HCR and Step, where lower is better.
5.3 Baseline
For comparison, we conducted controlled experiments with a BaseAgent, which ablates all three modules of VisEscaper on $exp_{\text{base}}$ setting. This experiment evaluated how much progress could be achieved solely with the recent history of actions and observations, excluding the constituent modules. For BaseAgent, only the most recent 20 observations and actions are alternately provided. We omitted open-source models since the success rate of BaseAgent with these models was 0 across all experiments.
According to Table 2, Claude-3.5-Sonnet, GPT-4o, and GPT-4o-mini all demonstrated significant improvements across all metrics-specifically increase $\times$3.7 in success rate, $\times$2.4 in goal completion, $\times$4.9 in SPL, and 7.3% decrease in total steps on average.
Notably, although higher success rates typically require more turns, all models with VisEscaper nevertheless demonstrated a reduction in the average step count, leading to SPL exhibiting the most substantial improvement among all metrics. This result underscores that the implementation of VisEscaper has made the exploration and decision-making processes of the original VLM agent both more effective and efficient.
5.4 Results
1. Even advanced models cannot achieve success without any hints. Table 3 presents the results of various VLMs integrated within the VisEscaper framework under both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Without any hints ($exp_{\text{base}}$), even Claude-3.5-Sonnet achieves only a 21.7% escape success rate and 32.37% goal completion, which is still 3-4 times lower than the human scores. Considering that humans show a minimal difference between the achievement rate and the final escape success rate (95.80 for GC and 82.5 for SR), the results from all models demonstrate that despite making some progress within each room, they cannot guarantee long-horizontal continuous success when no guidance is provided.
2. High-performing models show a drastic increase in all metrics, achieving objectives with fewer hints and within fewer turns. When hints are provided to stuck agents ($exp_{\text{hint}}$), there is a significant increase in success rate and goal completion across all models. Notably, Claude-3.5-Sonnet and GPT-4o escaped rooms using less than half the average number of steps compared to other models, making them much more efficient at path optimization, which leads to the greatest improvement, especially in SPL, among all metrics. Also, smaller models still improve substantially with hints, but rely on them more frequently, as shown by higher HCR.
3. High-performing models demonstrate accelerated goal completion rates in later stages compared to early stages. Figure 6 illustrates the progression of GC in Table 3 across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory.
Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) used fewer hints compared to other models (HCR of $exp_{\text{hint}}$ in Table 3). This indicates that rapid progress by high-performing models does not stem from heavy reliance on hints; instead, they use hints selectively only to resolve specific obstacles and, once past these obstacles, accelerate progress by effectively leveraging information accumulated up to that point.
Figure 6: Average GC (Goal Completion) progression across all rooms and experiments during the first 100 steps. The X-axis represents the number of steps, and the Y-axis represents the goal completion at each step. Solid lines represent results from the $exp_{\text{hint}}$ setting, while dotted lines represent those from the $exp_{\text{base}}$ setting.
6 Additional Analysis
To scrutinize the performance differences between various models in VisEscape, we conduct additional analysis on action repetition and impact of the Memory module.
$exp_{\text{base}}$ $exp_{\text{hint}}$
[…]
InternVL2.5-38B Image 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
InternVL2.5-38B + Base LLM (R1) Caption 10.3 19.58 7.81 261.6 80.0 87.06 25.53 19.37 166.8
InternVL2-40B Image 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-v1.6-34B + Base LLM Caption 0.0 4.14 0.0 - 38.3 61.80 50.08 4.59 257.7
LLaVA-OneVision-7B Image 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
LLaVA-OneVision-7B + Base LLM Caption 0.0 1.62 0.0 - 25.0 58.93 70.70 2.40 287.3
Table 4: Comparison of scores between different run types: using only VLM (which receives an image as input) and using a mixture of VLM with LLM (which receives a caption as input). Metrics improved with caption inputs are highlighted in bold. Base LLM refers to the language backbone underlying the corresponding VLM. Base LLM (R1) for InternVL2.5-38B denotes Deepseek-R1-Distill-Qwen-32B. Experiments without step counts all ended in failure due to reaching the maximum turn limit.
6.1 Repetitive Ineffective Action Patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table 3 despite the expectation that more steps would be required for higher SR and GC, we observe that lower-performing models frequently repeat the same action in identical situations throughout their trajectories. Figure 7 shows that these redundant attempts are more prevalent in lower-performing models, including both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Consequently, they accumulate longer trajectories, which in turn leads to lower SPL scores.
Figure 7: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions) on $exp_{\text{base}}$ setting. The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
[…]
InternVL2.5-38B O 66.7 79.2 41.7 12.2 210.5
X 10.0 59.2 41.9 1.6 287.0
LLaVA-v1.6-34B O 15.0 34.8 65.0 1.5 289.8
X 0.0 15.2 55.2 0.0 300.0
Table 5: Ablation results in $exp_{\text{hint}}$ comparing performance in the presence and absence of memory modules. The scores of all metrics decrease significantly when memory modules is absent.
6.2 Ablation of the Memory Module
The key characteristic of VisEscape —requiring the agent to retain and utilize information collected throughout the trajectory—highlights memory as essential for effectively navigating in partially observable environments. To investigate how and to what extent the Memory module contributes to VisEscaper, we conduct ablations on the Memory module, and analyze the degree of exploration for each condition using two models: InternVL2.5-38B and LLaVA-v1.6-34B. For the ablation, instead of using a Memory module, we enumerate all unique observations made up to the current step and feed them directly into the ReAct module. Table 5 shows a significant decrease across metrics due to absence of the Memory module. Notably, SR and SPL decreased more significantly than GC, indicating that without the Memory module, the agent struggles to retain the extensive information required to eventually escape the room.
6.3 Replacing VLM with LLM
Socratic models methodology, as represented by [36], is an approach that uses image captioners to convert visual inputs into descriptive text inputs and then employs LLMs to perform multimodal inference. Since it is widely known that VLMs excel at visual perception but often struggle with complex visual reasoning [22, 14, 26], we aim to address this gap by adopting LLMs.
Therefore, we compared the results between this paper’s default setting—where the VLM directly receives visual observations—and the Socratic Models methodology, where the LLM instead receives captions generated by the VLM. For controlled experiments, we used only the LLM that served as backbone language model of corresponding VLM  [^2] Refer to Appendix B.2. for the names of the corresponding LLMs for each VLM.. Results are presented in Table 4.
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed a decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substituting the LLM Qwen-32B-Instruct with Deepseek-R1-Distill-Qwen-32B, which shares an identical model architecture but has undergone reasoning capability distillation from a state-of-the-art reasoning model (Deepseek-R1 [15]) and thus possesses remarkably enhanced reasoning capabilities, all evaluation metrics improved notably. These results suggest that: 1. Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2. Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
7 Conclusion
In this study, we introduce VisEscape, a novel benchmark of escape room scenarios crafted to rigorously test multimodal agents’ abilities to effectively explore, reason, and make decisions in interactive and dynamic environments. We demonstrate that existing multimodal models struggle significantly under these demanding conditions. To address this issue, we propose VisEscaper framework, achieving substantial improvements in both effectiveness and efficiency. Our findings highlight the critical importance of memory management and adaptive reasoning capacity for multimodal agents, providing pathways toward developing more capable, autonomous AI systems for complex, real-world problem-solving tasks.
8 Limitation
While VisEscape presents a valuable benchmark for testing the exploration-driven problem-solving abilities of multimodal agents, the specific task of room escapes may not fully generalize to real-world environments where agents encounter more complex situations and more unfamiliar objects. Additionally, since the action space of VisEscape is discrete rather than continuous, there is a need to convert the conclusions drawn from multiple modules into feasible continuous action sequences for real-world execution.
9 Acknowledgements
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Development of multimodal UX evaluation platform technology for XR spatial responsive content optimization, Project Number: RS-2024-00361757).
Supplementary Material
Appendix Overview
  • In Appendix A, we provide results and analysis for additional experiments that were not covered in the main paper.
  • In Appendix B, we provide comprehensive details about our experimental setup and the prompt template used.
  • In Appendix C, we provide additional information about our data construction process.
  • In Appendix D, we present details about the conducted human study.
Appendix A. Additional Experiments and Analysis
A.1. Tendency of Brute-Force Approaches
In Figure 7, we show that low-performing agents exhibited meaningless, repetitive action patterns. However, upon analyzing the trajectories of various agents, we found that, when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempted various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated actions according to the following rules: (1) If the reasoning was grounded in meaningful clues or observations, we did not annotate it as brute-force. (2) If the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated them as brute-force.
Table 6 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL2.5-38B and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made a greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they made more attempts and fail more frequently on quizzes.
Model Brute-force attempt ratio(%) Solved quiz ratio(%)
Claude-3.5-Sonnet 0.0 21.2
GPT-4o 0.9 25.9
GPT-4o-mini 10.7 6.3
InternVL2.5-38B 10.5 6.6
LLaVA-v1.6-34B 70.0 0.0
Table 6: Results for puzzle-solving. Brute-force attempt ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape under the $exp_{\text{base}}$ setting.
A.2. Effect of the Exploration Memory
According to Table 5, removing the Memory module notably reduced VisEscaper performance. Beyond performance metrics, we hypothesize that among the three components of Memory module, particularly the Exploration memory component, which tracks already explored versus unexplored scenes or objects, encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the Exploration memory, while keeping the other two components (Structured spatial memory and Salient action memory) intact within the Memory module. For both settings, we measured the proportion of scenes observed by the agent from the set of essential observations (as determined by the oracle trajectory) required to successfully complete the game. Figure 9 shows that InternVL2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-v1.6-34B shows a substantial decrease.
A.3. Visualization of Interactions in VisEscape
Figure 8 is a visualization of interactions defined in VisEscape. The size of each word is proportional to its frequency of appearance within the game.
Figure 8: Wordcloud of all types of actions defined in VisEscape. We excluded “open” and “close”, which are available on most receptacles.
Figure 9: Proportion of essential scenes observed by the agent at least once within the initial 100 steps. Solid and dotted lines denote results with and without exploration memory, respectively. Removing exploration memory significantly reduces the number of observed essential scenes in both models.
A.4. Statistics for Image Captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the VLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information consistent with the GT state. Table 7 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise way, even when compared to GPT-4o.
Models Accuracy(%) Avg. Length
GPT-4o 73.9 47.03
GPT-4o-mini 72.8 64.46
LLaVA-v1.6-34B 63.8 91.19
InternVL2.5-38B 75.1 38.50
Table 7: Evaluation on captions generated by each VLMs. Accuracy indicates whether the caption accurately describes the current game state. Avg. Length denotes the average length of all captions generated by each model for every unique scene.
$exp_{\text{base}}$ $exp_{\text{hint}}$
Model(VLM / LLM) Input Modality SR(%) GC(%) SPL(%) Step SR(%) GC(%) HCR(%) SPL(%) Step
GPT-4o Image 13.3 29.02 3.91 276.2 95.0 97.39 20.74 26.83 112.9
GPT-4o Caption 10.0 23.54 3.42 280.7 98.3 98.00 17.39 33.60 91.8
GPT-4o-mini Image 6.7 21.31 1.05 291.2 46.7 67.40 32.16 8.93 228.2
GPT-4o-mini Caption 11.7 24.21 3.25 280.3 53.3 73.94 33.93 9.76 225.6
InternVL2.5-38B Image 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
InternVL2.5-38B + Base LLM (R1) Caption 10.3 19.58 7.81 261.6 80.0 87.06 25.53 19.37 166.8
InternVL2-40B Image 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-v1.6-34B + Base LLM Caption 0.0 4.14 0.0 - 38.3 61.80 50.08 4.59 257.7
LLaVA-OneVision-7B Image 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
LLaVA-OneVision-7B + Base LLM Caption 0.0 1.62 0.0 - 25.0 58.93 70.70 2.40 287.3
Table 8: Full results are shown in Table 4. For GPT-4o and GPT-4o-mini, we used the same model but varied the input modality type, since they do not provide language models that receives text input only.
A.5. Full Results of Table 4
Table 8 includes results for GPT-4o and GPT-4o-mini, which were omitted in Table 4. When GPT-4o-mini used captions instead of images as input, there was a slight improvement across all metrics, but it did not show a significant difference compared to other open-source models. Notably, GPT-4o’s performance decreased in $exp_{\text{base}}$ but slightly improved in $exp_{\text{hint}}$ setting.
Appendix B. Experimental details
B.1. List of LLMs and VLMs Used
In the list below, we list the versions of the LLMs used in our experiments. For GPT and Claude models, we use the officially released versions, and for the others, we use the model versions available on Huggingface.
  • Claude-3.5-Sonnet  [3]:
    claude-3-5-sonnet-20241022
  • GPT-4o  [1]:
    gpt-4o-2024-08-06
  • GPT-4o-mini  [1]:
    gpt-4o-mini
  • LLaVA-v1.6-34B [23]:
    llava-hf/llava-v1.6-34b-hf
  • LLaVA-OneVision-Qwen2-7B [20]:
    llava-hf/llava-onevision-qwen2-7b-ov-hf
  • InternVL2-40B [6]:
    OpenGVLab/InternVL2-40B
  • InternVL2.5-38B [6]:
    OpenGVLab/InternVL2_5-38B
  • Qwen2-7B-Instruct [42]:
    Qwen/Qwen2-7B-Instruct
  • Qwen2.5-32B-Instruct [43]:
    Qwen/Qwen2.5-32B-Instruct
  • Yi-34B [46]:
    NousResearch/Nous-Hermes-2-Yi-34B
  • DeepSeek-R1-Distill-Qwen-32B [15]:
    deepseek-ai/DeepSeek-R1-Distill-
    Qwen-32B
B.2. VLMs and Corresponding Base LLMs from Table 4
As shown in Table 4, we used only the LLM that served as backbone language model of corresponding VLM for controlled experiments. Base LLMs corresponding to each VLMS used in Table 4 is shown in Table 9333The documents specifying the LLMs selected as language backbone models are as follows: InternVL2.5-38B: https://huggingface.co/OpenGVLab/InternVL2_5-38B InternVL2-40B: https://huggingface.co/OpenGVLab/InternVL2-40B LLaVA-v1.6-34B: https://huggingface.co/llava-hf/llava-v1.6-34b-hf LLaVA-OneVision-7B: https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
Model Type Model Name
VLM InternVL2.5-38B
Base LLM Qwen2.5-32B-Instruct
Base LLM (R1) DeepSeek-R1-Distill-Qwen-32B
VLM InternVL2-40B
Base LLM Yi-34B
VLM LLaVA-v1.6-34B
Base LLM Yi-34B
VLM LLaVA-OneVision-7B
Base LLM Qwen2-7B-Instruct
Table 9: VLMs and their corresponding base LLMs used in experiments on Table 4.
B.3. Computational Resources and Environments
For closed-source models, we used the OpenAI API for running GPT-4o and GPT-4o-mini, and the Anthropic API for running Claude-3.5-Sonnet.
For open-source models, we downloaded models from Huggingface and served models using the vLLM server. We categorized the models as follows: small models (under 10B parameters), large models (30 to 40B parameters). For small models, we allocated a single NVIDIA RTX 4090 GPU per model. For large models, we allocated two or four NVIDIA A6000 GPUs per model.
All results reported in this paper are aggregated from three different runs conducted across 20 rooms. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often terminated early due to maximum turn limit) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 per room for three separate experiments, totaling about \$50 for a complete run.
B.4. Prompt templates
Table 1011121314 shows experiment prompts.
Variables:
direction, available_actions
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
Prompt:
{Initial Prompt}
<Current Observation>{direction} side of room - [IMAGE]</Current Observation>
Based on these information, choose next action to progress in the game. You can do one of the following actions: {available_actions}
Before you act, think first, and then act. Your thought should be in section [THINK], and your action should be in section [ACTION]. In [ACTION], respond ONLY with the chosen action, no other text.
[THINK]
[ACTION]
(a) Prompts for ReAct module on initial step
Variables:
memory, salient_action_history, action_history, direction, current_scene_desc, previous_react, available_actions, hint_guideline_text
Puzzle text [OPTIONAL: if ispuzzle mode]:
<ANSWER> is an action to input the answer to open the lock you are facing. When you choose <ANSWER>, you should follow this format: <̈ANSWER>your answer</ANSWER>.̈
Prompt (ReAct):
{Initial Prompt}
<Memory>{memory}</Memory>
<Action Memory>{salient_action_history}</Action Memory>
<Recent actions(from oldest to latest)>{action_history}</Recent actions>
<Current Observation>{direction} side of room - {current_scene_desc}</Current Observation>
<Your Thought and Action before this turn>{previous_react}</Your Thought and Action before this turn>
{available_actions}
{Puzzle text}
Before you act, think first, and then act. If there is a hint message, you should choose action to accomplish the guideline in hint message. {hint_guideline_text}
Your thought should be in section [THINK], and your action should be in section [ACTION]. In [ACTION], respond ONLY with the chosen action or <ANSWER>your answer</ANSWER>, no other text.
Prompt (Retry):
{Initial Prompt}
<Memory>{memory}</Memory>
<Your Previous Action> {before_action}
<Available Actions> {available_actions}
You just performed the <Your Previous Action>, but that action is not currently available in <Available Actions>. Referring to your memory, choose an action that is necessary to perform <Your Previous Action>.
{hint_guideline_text}
You should choose one of the following actions:{available_actions}
Please respond following below format without any other text: [ACTION]
(b) Prompts for ReAct module and Retry
Table 10: Prompts for the ReAct module, including the initial step and subsequent steps.
Variables:
action_history
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
last_10_history (for log in action_history) :
[…]
Action: [{log[’action’]}]-{log[’analysis’]}
spatial_json_format:
{"direction 1" : {
   "objects":["object1", "object2", ...]
inspected_objects_json_format:
[{"object 1" : {
   "state":"",
   "characteristics":"",
   "additional info":""
Prompt:
{Initial Prompt}
<Last 10 logs(from oldest to latest)> {last_10_history} </Last 10 logs>
Here is the definition of information:
<Last 10 logs>: Sequence of observation, action-effect, next observation, next action-effect for each turn.
Here is the task:
Construct your memory about the room, based on the last 10 logs. Follow below guidelines:
1. Identify which objects exist on each directional wall, and add the information to "[SPATIAL MEMORY]" section.
2. If you have inspected an object via "inspect [object]", you should add the information to "[INSPECTED OBJECTS]" section.
3. The objects that you have not inspected via "inspect [object]" should be added to "[UNINSPECTED OBJECTS]" section.
4. Add any information from observations which is not included in [SPATIAL MEMORY] and [INSPECTED OBJECTS] that you think is necessary for solving other problems to the "[Additional Memory]" section.
Please respond following below format without any other text:
[SPATIAL MEMORY] {spatial_json_format}
[INSPECTED OBJECTS] {inspected_objects_json_format}
[UNINSPECTED OBJECTS] []
[ADDITIONAL MEMORY] [1. additional memory1, 2. additional memory2, ...]
Table 11: Prompts for Memory module - for first run (memory construction).
Variables:
spatial_memory, action_history
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
last_10_history (for log in action_history) :
[…]
Action: [{log[’action’]}]-{log[’analysis’]}
spatial_json_format:
{"direction 1" : {
   "objects":["object1", "object2", ...]
inspected_objects_json_format:
[{"object 1" : {
   "state":"",
   "characteristics":"",
   "additional info":""
Prompt:
{Initial Prompt}
<Current Memory> {spatial_memory} </Current Memory>
<Last 10 logs(from oldest to latest)> {last_10_history} </Last 10 logs>
Here is the definition of each information:
<Last 10 logs>: Sequence of observation, action-effect, next observation, next action-effect for each turn.
Here is the task:
Update your memory about the room, based on the last 10 logs. Follow below guidelines:
1. If you newly inspected an object via "inspect [object]" among objects in "[UNINSPECTED OBJECTS]" section,
  you should add the information to "[INSPECTED OBJECTS]" section.
2. Based on the information that you can obtain from last 10 logs, update <Current Memory>
  if there exists any information that you can obtain from <Last 10 logs> but not in <Current Memory>.
3. Add any information not included in spatial memory and inspected objects that you think is necessary
  for solving other problems to the "[ADDITIONAL MEMORY]" section.
Please respond following below format without any other text:
[SPATIAL MEMORY] {spatial_json_format}
[INSPECTED OBJECTS] {inspected_objects_json_format}
[UNINSPECTED OBJECTS] []
[ADDITIONAL MEMORY] [1. additional memory1, 2. additional memory2, ...]
Table 12: Prompts for Memory module (memory update).
Variables:
previous_scene_desc, previous_action, current_scene_desc
Prompt:
{Initial Prompt}
<Previous Observation> : {previous_scene_desc}
<Previous Action> : {previous_action}
<Current Observation> : {current_scene_desc}
Here is the definition of each information:
<Previous Observation> is a description of a scene before your action, and <Current Observation> is a description of a scene after your action.
Here is the task:
Analyze the effect of your action by comparing <Previous Observation> and <Current Observation>. The analysis should be concise and definitive, not descriptive.
Keep it under 10 words.
[ANALYSIS]
Table 13: Prompts for Feedback module.
Variables:
item_name
Prompt:
This image is a close-up view of an item ’{item_name}’.
Describe this image. Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(a) Prompts for getting caption from image of item-view
Variables:
object_type, items_str
This image is a close-up view of an object ’{object_type}’.
In {object_type}, the following objects are present: {items_str}
Describe this image. The names of visible objects should be expressed using the given object names above, enclosed in "".
Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not include any analysis of the scene or the room, just describe the image.
3. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(b) Prompts for getting caption from image of object-view
Variables:
objects
This image is a wall view of a room, with following objects: {objects}.
Describe this image. The names of visible objects should be expressed using the given object names above, enclosed in "".
Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not include any analysis of the scene or the room, just describe the image.
3. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(c) Prompts for getting caption from image of wall-view
Table 14: Prompts for captioning.
Appendix C. Dataset Construction Process
C.1. Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined states, such as open/closed, locked/unlocked, and pushed/pulled, enabling agent interactions. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items are shown in in Figure 14.
C.2. Scene Creation Using Autodesk Revit
We employ Autodesk Revit to construct room escape scenes by integrating the assets created in Trimble SketchUp. Autodesk Revit enables the precise placement of receptacles and objects within the scene to simulate realistic escape room environments. Each scene includes multiple viewpoints, which are categorized into three types:
  • Wall View: A view showing an entire wall on one side of east, west, north, or south. This view updates dynamically according to the states of receptacles and items belonging to each wall.
  • Receptacle View: A close-up view of an receptacle accessed through the “inspect” action from the Wall View. It changes to reflect the state of receptacles.
  • Item View: A close-up view of an item accessed through the “inspect” action from the Receptacle View. It provides a close-up perspective of critical elements such as locks, puzzles, and other objects directly involved in problem solving.
Figure 14: Objects created using Trimble Sketchup for VisEscape. On the left are objects of the Receptacle type, and on the right are objects of the Item type.
Figure 15: Before and after applying visual rendering using Chaos Enscape. The left shows the view in Autodesk Revit before using Chaos Enscape, and the right shows the view in Chaos Enscape.
C.3. Enhancing Photo-realism Using Chaos Enscape
To enhance the realism of the dataset, we use Chaos Enscape, a rendering tool that processes scene data generated in Autodesk Revit. Chaos Enscape improves the visual fidelity of the scene, providing photorealistic object representations. This increased realism facilitates effective agent reasoning and interaction within the environment, enabling accurate simulation and analysis of escape room scenarios. Refer to Figure 15 to check the before and after application of Chaos Enscape,
Figure 16: Example of rooms rendered using Chaos Enscape.
Appendix D. Human Study
Room SR(%) GC(%) Steps Duration(s)
1 75.0 92.5 41.8 231.3
2 100.0 100.0 67.3 319.5
3 50.0 87.5 71.1 319.5
4 75.0 93.8 112.0 447.0
5 100.0 100.0 77.3 274.5
6 100.0 100.0 37.0 117.0
7 100.0 100.0 46.8 225.0
8 100.0 100.0 53.0 567.3
9 75.0 90.0 52.0 567.3
10 75.0 91.7 54.0 204.5
11 50.0 79.2 37.8 231.3
12 75.0 92.9 31.3 144.0
13 100.0 100.0 41.5 199.8
14 100.0 100.0 57.3 468.3
15 100.0 100.0 36.0 151.0
16 100.0 100.0 40.8 106.0
17 25.0 96.9 63.3 488.3
18 100.0 100.0 40.5 399.3
19 75.0 91.7 43.8 378.5
20 75.0 100.0 50.8 530.3
Average 82.5 95.8 52.8 318.5
Table 15: Human evaluation per-room success rates, goal completions, step counts, and time durations.
For the human study, Figures 17 and 18 illustrate the Gradio interface used to implement the game environment.
Table 15 presents the results of the human evaluation for 20 rooms, including the success rate, goal completion, number of steps, and time duration for experiments. Participants achieved a high success rates in most rooms. We recruited 20 participants for this evaluation and ensured fair compensation according to standard ethical guidelines. Each participant evaluated 4 rooms.
Figure 17: Human evaluation using the Gradio interface.
Figure 18: Human evaluation using the Gradio interface.