### v3

# Instructions for *ACL Proceedings

Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observe that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 4.9 times more efficiently on average compared to baseline agents.
Instructions for *ACL Proceedings
Anonymous ACL submission
1 Introduction
Realistic goals often necessitate exploration; for pantry cooking, one must search through cabinets to identify available ingredients and determine how they can be combined into a meal. Likewise, a trip to a foreign city cannot be planned without searching for major sights and train tables. In such cases, effective planning depends on acquiring information through interaction with the environment rather than having it provided a priori.
Existing works on embodied agents have addressed the importance of exploration, but it has primarily been studied under decomposable and explicit tasks like “Place a cleaned egg in a microwave” Shridhar_2020_CVPR; puig2018virtualhome; li2021igibson; kim2024realfred. However, the capability of agents to autonomously search for and utilize relevant information from environments where tasks or solutions are not explicitly defined has not been sufficiently investigated.
Figure 1: In VisEscape, player actions cause state changes that may or may not be immediately visible. For instance, pressing a button might move an object located outside the player’s immediate field of view. These key characteristics—partial observability and dynamic state changes—require players to maintain spatial awareness and continually update their memory.
Figure 2: An illustration of an excerpt of a trajectory from VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes in different locations to solve creative puzzles.
To bridge this gap, we argue that practical evaluation of decision-making agents should assess their ability to conduct self-directed exploration to identify underlying objectives and subsequently to take reasoned actions. In particular, this entails formulating hypotheses for problem-solving based on information and clues gathered through exploration.
To evaluate agents under these conditions, we propose VisEscape, a benchmark of 20 virtual escape rooms for assessing exploration-centric problem-solving capacities of multimodal agents. Given the implicit and unspecified goal of “escape the room”, the agents explore their surroundings, identify and integrate relevant information, discover how objects can be used, and solve complex puzzles to successfully escape. Moreover, they need to retain previously observed information and interpret indirect clues, requiring them to reason beyond immediate observations and dynamically adapt to environmental changes.
On VisEscape, we evaluate various VLMs’ performance and observe that even advanced models like Claude-3.5-Sonnet and GPT-4o show low escape rates of only 10%, while all other models achieve 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Reasoning, and Reflection—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escaping successfully escaped, while models with pre-existing escape capabilities demonstrate $\times$3.7 enhanced effectiveness and $\times$4.9 improved efficiency. Additionally, [TODO: findings from analysis experiment]
In summary, the key points of this paper are threefold:
  • We introduce VisEscape, a unique and challenging benchmark for evaluating exploration-centric problem-solving abilities of multimodal agents in dynamic environments.
  • We propose VisEscaper framework, which significantly improves multimodal agents’ performance in terms of effectiveness and efficiency.
  • We address the deficiencies in reasoning capacity of VLMs by leveraging the reasoning capacity of LLMs, thereby offering valuable insights into potential improvements for multimodal agents.
2 Key Challenges of VisEscape
Figure 2 illustrates an example of the problem-solving process required in the VisEscape. The game agent should execute a sequence of tasks comprising diverse interactions.
1. Self-directed Exploration. As in real escape room games, the game agent in VisEscape receives no instruction except for the instruction ’escape the room’. Therefore, without a clear roadmap, the agent must actively explore and figure out how to proceed through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement.
2. Long-Horizontal successive task. Under the uncertainty about how to progress without explicit guidance, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires an average of 27.2 steps to successfully escape, even for the oracle trajectory. The considerable length of the task sequence and the dependencies between checkpoints require agents to effectively leverage information from past trajectory. Success hinges on the ability to integrate these prior discoveries into the ongoing decision-making process.
3 Details on Dataset Composition and Construction
3.1 Dataset Composition
Title. VisEscape consists of 20 distinct rooms. Each room comprises four walls corresponding to four directions, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in Shridhar_2020_CVPR. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room contains an average of 5.9 receptacles and 5.1 items, along with one or two visual quizzes. Statistics and examples are shown in Table 1 and Figure 14LABEL:fig:example.
Game Graph. To track the dynamically changing room state and provide corresponding observations, we used a Finite State Machine (FSM). FSM consists of states and transitions-we mapped FSM states to observations matching the current game state and modeled how agent’s actions change observations through FSM state transitions. See Appendix LABEL: for more details.
Scene Receptacle Item Quiz Interaction
# Total 1084 118 102 27 311
# Types 662 32 23 27 37
Avg./room 54.2 5.9 5.1 1.4 15.6
Table 1: Statistics for scenes, receptacles, items, quizzes, and interactions in VisEscape.
Actions. Agent can perform two types of actions: movement and object interaction. For movement, it can “inspect [object]” to examine an object closely or step back to gain a broader view. Additionally, it can use turn to [direction] to view other walls that are not currently observable. After moving, a corresponding observation is given based on the node connections in the game graph. For interaction with objects, the agent can utilize acquired items (e.g., Unlock with key) or perform actions that are set for the objects it can see (§2.2). When an agent performs an action that satisfies the object’s predefined conditions, it triggers state transitions according to game graph. See Appendix LABEL: for examples of diverse object interactions.
Figure 3: Process of VisEscape construction.
3.2 VisEscape Construction
Figure 3 shows the creation process for rooms in VisEscape. (a) Objects are created by 3D modeling software, and those needed for arrangement in each room were sampled. (b) Using the sampled objects, game logic defining the key checkpoints to clear the game was designed with AI-human collaboration. A game graph is then built as a FSM to track the overall game state and enable state transitions triggered by agent’s actions. (c) Objects were arranged in the space according to the room’s setup, and visual observations corresponding to game states were mapped. See Appendix LABEL: for detailed explanations of each process, and the prompts used for game logic design.
Frontier MLLMs Open-source MLLMs
[…]
Human 82.5 95.80 51.30 52.8
Table 2: Performance comparison of various MLLMs and human performance on the VisEscape benchmark. The best results for each metric are bolded. Refer Appendix G for detailed information on the human study. For models that failed to escape entirely, the standard deviation is marked with "-".
4 Experiments
4.1 Experiment Setup
At each step, the model receives the current game state observation, the history of the last 20 observations and actions, held items, and available actions, from which it selects one executable action. For each room, we conducted three different trials and averaged the resulting metrics. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded after a total of 300 steps.
4.2 Metrics
We adopted the major metrics widely used in benchmarks for embodied agents. These include: 1. Success Rate (SR) , which is the ratio of escape success in an episode; 2. Goal Completion (GC) , the ratio of checkpoints completed at the end of an episode; 3. Step, the total actions taken by an agent; and 4. Success weighted by Path Length (SPL) anderson2018evaluation, which is the success rate weighted by the term:
Path length of the oracle trajectoryPath length taken by the $\frac{\text{Path length of the oracle trajectory}}{\text{Path length taken by the agent}}$ (1)
This term penalizes the success rate based on the total number of steps taken.
For $exp_{\text{hint}}$, we define a metric regarding the agent’s hint usage: Hint-assisted Completion Rate (HCR), which denotes the ratio of checkpoints completed by the agent with the assistance of hints.
4.3 Results
Table 2 presents the experimental results for a range of closed-source and open-source VLMs. Without any hints, None of the open-source models achieved a single success. Notably, even sophisticated models such as Claude 3.5 Sonnet and GPT-4o demonstrated success rates below 10 percent. These findings underscore the highly challenging nature of the exploration-centric problem-solving environment provided by VisEscape for current models, concurrently highlighting significant avenues for improvement.
To improve model proficiency in VisEscape, agents must effectively execute two primary capabilities that VisEscape requires: exploration and iterative hypothesis formulation and verification. Rather than devising a specialized framework for the unique ’escape room’ task, we aim to analyze how modules considered crucial in agent modeling contribute to the exploration-centric problem-solving task provided by this benchmark. To this end, we propose VisEscaper, augmented with Memory, Self-feedback, and Reasoning modules, and conduct studies for analysis.
5 Can Memory Management helps Exploration?
In VisEscape, agent can only observe a limited portion of the room-for effective navigation and exploration, the agent should maintain spatial memory that the agent constructs for itself so that it can form a coherent mental map of the space. Additionally, it must remember the actions it has attempted so that it can learn from its trials and avoid redundant efforts.
5.1 Memory module Design
Therefore, we designed and applied a Memory Module to the game agent. The memory module consists of Structured Spatial Memory, Exploration Memory, and Salient Action Memory (Refer to Figure 4), which are constructed and updated periodically. It compressively and structurally manages memory from observation and action histories accumulated over tens to hundreds of steps, while also updating periodically. This allows the agent to make clear decisions from a condensed history, effectively mitigating its processing overload by addressing both cognitive demands and the model’s inherent context window constraints.
5.2 Memory and Effective Exploration
Figure 4: [TODO: how about iteration in figure?]
6 How Game Agents perform reasoning?
The Reasoning module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research yao2023react, the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition, the agent transmits its reasoning from the previous step to the next step, enabling it to maintain consistent short-term plans across multiple steps.
6.1 Ablation on Reasoning
To analyze how performing reasoning prior to action decisions affects 1) the action decisions themselves and 2) game performance, we conducted an ablation study on the reasoning module.
6.2 Performance of VisEscaper
Model $A_{R}$= $A$ $A_{R}$$\neq$ $A$
$A_{R}$> $A$ $A_{R}$< $A$ ambig.
Claude 3.5 Sonnet 0.47 0.5 0.1 0.4
GPT-4o 0.54 0.1 0.4 0.4
Gemini-1.5-Pro 0.52 0.5 0.1 0.4
GPT-4o-mini 0.53 0.1 0.4 0.4
Table 3: Generated by Spread-LaTeX
7 Memory management helps effective exploration.
7.1 Memory
Memory helps effective exploration.
8 Additional Analysis
To scrutinize the performance differences between various models in VisEscape, we conduct additional analysis on action repetition and impact of the Memory module.
$exp_{\text{base}}$ $exp_{\text{hint}}$
[…]
InternVL2.5-38B Image 0.0 6.20 - - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
InternVL2.5-38B + Base LLM (R1) Caption 10.3 19.58 7.81 261.6 80.0 87.06 25.53 19.37 166.8
InternVL2-40B Image 0.0 4.91 - - 8.3 48.18 61.61 0.96 291.1
InternVL2-40B + Base LLM Caption 0.0 3.88 - - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 - - 15.0 34.79 64.97 1.47 289.8
LLaVA-v1.6-34B + Base LLM Caption 0.0 4.14 - - 38.3 61.80 50.08 4.59 257.7
LLaVA-OneVision-7B Image 0.0 1.25 - - 8.3 41.02 62.89 0.80 294.9
LLaVA-OneVision-7B + Base LLM Caption 0.0 1.62 - - 25.0 58.93 70.70 2.40 287.3
Table 4: Comparative performance of results from MLLMs and Socratic method. Experiment is conducted by models with memory management and reasoning module. $exp_{\text{base}}$ and $exp_{\text{hint}}$ denote experiments conducted without hints and hint-guided experiments, respectively.
8.1 Repetitive Ineffective Action Patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table LABEL:tab:main despite the expectation that more steps would be required for higher SR and GC, we observe that lower-performing models frequently repeat the same action in identical situations throughout their trajectories. Figure 10 shows that these redundant attempts are more prevalent in lower-performing models, including both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Consequently, they accumulate longer trajectories, which in turn leads to lower SPL scores.
Figure 5: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions). The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
[…]
InternVL2.5-38B O 66.7 79.2 41.7 12.2 210.5
X 10.0 59.2 41.9 1.6 287.0
LLaVA-v1.6-34B O 15.0 34.8 65.0 1.5 289.8
X 0.0 15.2 55.2 0.0 300.0
Table 5: Ablation results in $exp_{\text{hint}}$ comparing performance in the presence and absence of memory modules. The scores of all metrics decrease significantly when memory modules is absent.
8.2 Ablation of the Memory Module
The key characteristic of VisEscape —requiring the agent to retain and utilize information collected throughout the trajectory—highlights memory as essential for effectively navigating in partially observable environments. To investigate how and to what extent the Memory module contributes to VisEscaper, we conduct ablations on the Memory module, and analyze the degree of exploration for each condition using two models: InternVL2.5-38B and LLaVA-v1.6-34B. For the ablation, instead of using a Memory module, we enumerate all unique observations made up to the current step and feed them directly into the ReAct module. Table 5 shows a significant decrease across metrics due to absence of the Memory module. Notably, SR and SPL decreased more significantly than GC, indicating that without the Memory module, the agent struggles to retain the extensive information required to eventually escape the room.
8.3 Replacing VLM with LLM
Socratic models methodology, as represented by shin2024socratic, is an approach that uses image captioners to convert visual inputs into descriptive text inputs and then employs LLMs to perform multimodal inference. Since it is widely known that VLMs excel at visual perception but often struggle with complex visual reasoning li2024enhancing; gan2022vision; liu2023matcr, we aim to address this gap by adopting LLMs.
Therefore, we compared the results between this paper’s default setting—where the VLM directly receives visual observations—and the Socratic Models methodology, where the LLM instead receives captions generated by the VLM. For controlled experiments, we used only the LLM that served as backbone language model of corresponding VLM  [^1] Refer to Appendix B.2. for the names of the corresponding LLMs for each VLM.. Results are presented in Table 4.
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed a decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substituting the LLM Qwen-32B-Instruct with Deepseek-R1-Distill-Qwen-32B, which shares an identical model architecture but has undergone reasoning capability distillation from a state-of-the-art reasoning model (Deepseek-R1 guo2025deepseek) and thus possesses remarkably enhanced reasoning capabilities, all evaluation metrics improved notably. These results suggest that: 1. Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2. Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
9 Conclusion
In this study, we introduce VisEscape, a novel benchmark of escape room scenarios crafted to rigorously test multimodal agents’ abilities to effectively explore, reason, and make decisions in interactive and dynamic environments. We demonstrate that existing multimodal models struggle significantly under these demanding conditions. To address this issue, we propose VisEscaper framework, achieving substantial improvements in both effectiveness and efficiency. Our findings highlight the critical importance of memory management and adaptive reasoning capacity for multimodal agents, providing pathways toward developing more capable, autonomous AI systems for complex, real-world problem-solving tasks.
10 Related Works
10.1 Interactive Environments for Evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world cote2019textworld; hausknecht2020interactive; qian2024escapebench. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge wang2022scienceworld and ethical decision-making pan2023rewards.
Several studies have advanced LLM/VLM-based agents built on Minecraft wangvoyager; Odyssey2024; dong2024villageragent; guss2019minerl. In parallel, embodied agent simulators have facilitated evaluating language-driven agents interacting within visually grounded environments ai2thor; puig2018virtualhome; shridharalfworld; habitat19iccv.
Meanwhile, the GUI environment has recently been proposed as a benchmark deng2024multi; deng2023mind2web; xu2021grounding for evaluating VLMs’ interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfacesliuvisualwebbench; cheng2024seeclick, reason about their functionality shahbandeh2024naviqate, and execute actions to successfully complete given tasks shi2017world; furuta2023exposing; yao2022webshop; zhouwebarena.
10.2 LLM/VLM as an Action Decision Maker
Deploying LLM or VLM agents in real-world applications requires effective handling of dynamic and unpredictable environments mnih2013playing. Such environments are characterized by changing contexts, unexpected events, and continuous novel observations nagabandi2019learning; zhang2018adaptive. To successfully adapt, agents must actively self-evolve using mechanisms for adaptive reasoning and continuous learning zhou2024symbolic.
Recent studies have highlighted three critical capabilities for agent adaptation. First, accumulation of experience enables agents to remember previous interactions, particularly failures, and use these experiences to improve future decisions feng2024agile. Second, strategic exploration allows agents to intelligently choose actions based on past outcomes and external feedback, balancing the benefits of learning against potential risks zhu2023ghost. Third, knowledge abstraction helps agents form generalized high-level representations from raw observations, significantly improving their adaptability in diverse situations zhao2023expel.
To support these adaptation skills, we integrate three specialized modules into our agent architecture; Memory, Feedback, and ReAct yao2023react. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
Appendix A Example Appendix
Appendix Overview
  • In Appendix B, we provide details about our experimental setup and the prompt template used.
  • In Appendix C, we provide additional information about our data construction process.
  • In Appendix D, we provide results and analysis for additional experiments that were not covered in the main paper.
  • In Appendix  E, we provide additional details not extensively covered in the main paper.
  • In Appendix F, we briefly introduce and analyze works that address the topic of escape rooms.
  • In Appendix G, we present details about the human study conducted.
Appendix B Experimental details
B.1 Model List
In the list below, we list the versions of the LLMs used in our experiments. For GPT, Gemini, and Claude models, we use the officially released APIs, and for the others, we use the model versions available on Huggingface.
  • Claude-3.5-Sonnet  claude3:
    claude-3-5-sonnet-20241022
  • GPT-4o  achiam2023gpt:
    gpt-4o-2024-08-06
  • GPT-4o-mini  achiam2023gpt:
    gpt-4o-mini
  • gemini-1.5-Pro  team2024gemini:
    Gemini-1.5-Pro
  • gemini-2.0-Flash  google2024gemini:
    Gemini-2.0-Flash
  • Qwen2.5-VL-32B-Instruct bai2025qwen2_5vl:
    Qwen/Qwen2.5-VL-32B-Instruct
  • LLaVA-v1.6-34B liu2023visual:
    llava-hf/llava-v1.6-34b-hf
  • LLaVA-OneVision-Qwen2-7B li2024llava:
    llava-hf/llava-onevision-qwen2-7b-ov-hf
  • InternVL2-40B chen2024expanding:
    OpenGVLab/InternVL2-40B
  • InternVL2.5-38B chen2024expanding:
    OpenGVLab/InternVL2_5-38B
  • Qwen2-7B-Instruct yang2024qwen2:
    Qwen/Qwen2-7B-Instruct
  • Qwen2.5-32B-Instruct yang2024qwen2_5:
    Qwen/Qwen2.5-32B-Instruct
  • Yi-34B young2024yi:
    NousResearch/Nous-Hermes-2-Yi-34B
  • DeepSeek-R1-Distill-Qwen-32B guo2025deepseek:
    deepseek-ai/DeepSeek-R1-Distill-
    Qwen-32B
B.2 MLLMs and Corresponding Backbone LLMs
As shown in Table 4, we used only the LLM that served as backbone language model of corresponding MLLM for controlled experiments. Base LLMs corresponding to each MLLMs used in Table 4 are shown in Table 6222The documents specifying the LLMs selected as language backbone models are as follows: InternVL2.5-38B:
https://huggingface.co/OpenGVLab/InternVL2_5-38B
InternVL2-40B:
https://huggingface.co/OpenGVLab/InternVL2-40B
LLaVA-v1.6-34B:
https://huggingface.co/llava-hf/llava-v1.6-34b-hf
LLaVA-OneVision-7B:
https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
Model Type Model Name
VLM InternVL2.5-38B
Base LLM Qwen2.5-32B-Instruct
Base LLM (R1) DeepSeek-R1-Distill-Qwen-32B
VLM InternVL2-40B
Base LLM Yi-34B
VLM LLaVA-v1.6-34B
Base LLM Yi-34B
VLM LLaVA-OneVision-7B
Base LLM Qwen2-7B-Instruct
Table 6: MLLMs and their corresponding base LLMs.
B.3 Computational Resources and Environments
For closed-source models, we used the OpenAI API for running GPT-4o and GPT-4o-mini, and the Anthropic API for running Claude-3.5-Sonnet.
For open-source models, we downloaded models from Huggingface and served models using the vLLM server. We categorized the models as follows: small models (under 10B parameters), large models (30 to 40B parameters). For small models, we allocated a single NVIDIA RTX 4090 GPU per model. For large models, we allocated two or four NVIDIA A6000 GPUs per model.
All results reported in this paper are aggregated from three different runs conducted in 20 rooms. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often terminated early due to the maximum turn limit) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 per room for three separate experiments, totaling about \$50 for a complete run.
B.4 Hints
To facilitate the evaluation of the agent’s progression along its path, all game logic in VisEscape is designed to be sequential. Therefore, the next checkpoint the agent must achieve after its current one is predetermined, as is the hint to guide it there. Hint messages are provided only during Hint-guided experiments if the agent gets stuck for 30 or more steps (i.e., shows no change in checkpoint progression).
Here are two examples of hint messages.
  • Object interaction: “Use key to unlock the wardrobe at the south wall.”
  • Answer for lock: “Solve the numberlock in the safe at the east wall. The answer is 8056.”
B.5 Prompt templates
Table 1415161718 shows experiment prompts.
Appendix C Dataset Construction Process
C.1 Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined states, such as open/closed, locked/unlocked, and pushed/pulled, enabling agent interactions. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items are shown in Figure 6.
Figure 6: Objects created using Trimble Sketchup for VisEscape. On the left are objects of the Receptacle type, and on the right are objects of the Item type.
Figure 7: Comparison of applying visual rendering using Chaos Enscape. The left shows the view in Autodesk Revit before using Chaos Enscape, and the right shows the view in Chaos Enscape.
C.2 Scene Creation Using Autodesk Revit
We employ Autodesk Revit to construct room escape scenes by integrating the assets created in Trimble SketchUp. Autodesk Revit enables the precise placement of receptacles and objects within the scene to simulate realistic escape room environments. Each scene includes multiple viewpoints, which are categorized into three types:
  • Wall View: A view showing an entire wall on one side of east, west, north, or south. This view updates dynamically according to the states of receptacles and items belonging to each wall.
  • Receptacle View: A close-up view of a receptacle accessed through the “inspect” action from the Wall View. It changes to reflect the states of receptacles.
  • Item View: A close-up view of an item accessed through the “inspect” action from the Receptacle View. It provides a close-up perspective of critical elements such as locks, puzzles, and other objects directly involved in problem solving.
C.3 Enhancing Photo-realism Using Chaos Enscape
To enhance the realism of the dataset, we use Chaos Enscape, a rendering tool that processes scene data generated in Autodesk Revit. Chaos Enscape improves the visual fidelity of the scene, providing photorealistic object representations. This increased realism facilitates effective agent reasoning and interaction within the environment, enabling accurate simulation and analysis of escape room scenarios. Refer to Figure 7 to check the before and after application of Chaos Enscape. Visualization of other rooms is in Figure 22.
C.4 Game Logic Design
We used OpenAI’s web interface GPT models to design efficient and automated game logic. The authors then manually reviewed the generated game logic design to ensure error-free gameplay. The prompt used for game logic design is in Table 19.
Appendix D Additional Experiments and Analysis
D.1 Tendency of Brute-Force Approaches
We found that when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempted various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated actions according to the following rules: (1) If the reasoning was grounded in meaningful clues or observations, we did not annotate it as brute-force. (2) If the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated them as brute-force.
Table 7 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL2.5-38B, and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made a greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they made more attempts and failed more frequently on quizzes.
Model Brute-force attempt ratio(%) Solved quiz ratio(%)
Claude-3.5-Sonnet 0.0 21.2
GPT-4o 0.9 25.9
GPT-4o-mini 10.7 6.3
InternVL2.5-38B 10.5 6.6
LLaVA-v1.6-34B 70.0 0.0
Table 7: Results for puzzle-solving. Brute-force attempt ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape.
D.2 Reasoning and Visual quizzes
In VisEscape, all visual quizzes require associative thinking abilities. Analyzing game trajectories of diverse models, we find that most failures were on visual quizzes. So, we investigated whether these were problems they inherently couldn’t solve, or if they failed because they couldn’t properly associate the key clues within the escape room environment. For this, we reformed them as multi-image Visual Question Answering (VQA) problems. By providing the model with both the clue image and the target lock image, we inform them that the two images are related to each other, and then have them solve the problem. Used prompts are in Table 20.
Furthermore, we compare pass@1 and pass@10 to determine if models could correctly solve the task via iterative reasoning, even when their initial attempt failed. To support this repeated hypothesizing and testing, the model’s previously generated reasoning sentence for an incorrect answer, along with the incorrect answer itself, was fed back as input for the subsequent attempt.
We selected one visual quiz from each room and then evaluated the performance of each model. Table 8 shows the number of correct answers by each model on these 20 visual quizzes. Results show that open-source models generally demonstrate lower performance (Pass@1 without reasoning) on these visual quizzes compared to frontier models. Also, Reasoning tends to improve the number of correctly solved quizzes across all models.
While frontier models like GPT-4o and Claude-3.5-Sonnet show substantial performance gains with reasoning, the extent of this improvement varies, with some open-source models also showing notable gains, particularly when more attempts are allowed (e.g., InternVL2.5-38B shows a +5 improvement in Pass@10 with reasoning).
Also, iterative reasoning, as indicated by the general increase in scores from Pass@1 to Pass@10 under reasoning conditions, leads to a higher number of correct answers. This suggests that an iterative process of hypothesis formulation and testing enables the models to reach the correct solutions.
Model Pass@1 Pass@10
No Reason Reason No Reason Reason
[…]
Table 8: Comparison of pass@1 and pass@10 scores for models: performance with and without reasoning.
D.3 Effect of the Exploration Memory
We hypothesize that among the three components of memory management module, particularly the Exploration memory component, which tracks already explored versus unexplored scenes or objects, encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the Exploration memory, while keeping the other two components (Structured spatial memory and Salient action memory) intact within the Memory module. For both settings, we measured the proportion of scenes observed by the agent from the set of essential observations (as determined by the oracle trajectory) required to successfully complete the game. Figure 8 shows that InternVL2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-v1.6-34B shows a substantial decrease.
Figure 8: Proportion of essential scenes observed by the agent at least once within the initial 100 steps. Solid and dotted lines denote results with and without exploration memory, respectively. Removing exploration memory significantly reduces the number of observed essential scenes in both models.
D.4 Comparison on Progress
Figure 9 illustrates the progression of goal completion across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory.
Figure 9: Average GC (Goal Completion) progression across all rooms and experiments during the first 100 steps. The X-axis represents the number of steps, and the Y-axis represents the goal completion at each step. Solid lines represent results from hint-guided experiment, while dotted lines represent those without any hints.
Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) used fewer hints compared to other models (low HCR in Table LABEL:tab:main_hint). This indicates that rapid progress by high-performing models does not stem from heavy reliance on hints; instead, they use hints selectively only to resolve specific obstacles and, once past these obstacles, accelerate progress by effectively leveraging information accumulated up to that point.
D.5 Repetition
To analyze the performance difference between frontier models and open-source models, we analyzed repetitive actions (repeating the same action in identical situations) exhibited by each model within their trajectories. We observe that lower-performing models frequently repeat actions they have already taken throughout their trajectories, getting stuck in repetitive cycles. Figure 10 shows that these redundant attempts are more prevalent in lower-performing models.
Figure 10: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions). The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
D.6 Statistics for Image Captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the MLLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information consistent with the GT state. Table 9 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise way, even when compared to GPT-4o.
Model Accuracy(%) Avg. Length
GPT-4o 73.9 47.03
GPT-4o-mini 72.8 64.46
LLaVA-v1.6-34B 63.8 91.19
InternVL2.5-38B 75.1 38.50
Table 9: Evaluation on captions generated by each VLMs. Accuracy indicates whether the caption accurately describes the current game state. Avg. Length denotes the average length of all captions generated by each model for every unique scene.
$exp_{\text{base}}$ $exp_{\text{hint}}$
Model(VLM / LLM) Input Modality SR(%) GC(%) SPL(%) Step SR(%) GC(%) HCR(%) SPL(%) Step
GPT-4o Image 13.3 29.02 3.91 276.2 95.0 97.39 20.74 26.83 112.9
GPT-4o Caption 10.0 23.54 3.42 280.7 98.3 98.00 17.39 33.60 91.8
GPT-4o-mini Image 6.7 21.31 1.05 291.2 46.7 67.40 32.16 8.93 228.2
GPT-4o-mini Caption 11.7 24.21 3.25 280.3 53.3 73.94 33.93 9.76 225.6
InternVL2.5-38B Image 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
InternVL2.5-38B + Base LLM (R1) Caption 10.3 19.58 7.81 261.6 80.0 87.06 25.53 19.37 166.8
InternVL2-40B Image 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-v1.6-34B + Base LLM Caption 0.0 4.14 0.0 - 38.3 61.80 50.08 4.59 257.7
LLaVA-OneVision-7B Image 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
LLaVA-OneVision-7B + Base LLM Caption 0.0 1.62 0.0 - 25.0 58.93 70.70 2.40 287.3
Table 10: Full results are shown in Table 4. For GPT-4o and GPT-4o-mini, we used the same model but varied the input modality type, since they do not provide language models that receives text input only.
D.7 Full Results of LLM Experiment
Table 10 includes results for GPT-4o and GPT-4o-mini, which were omitted in Table 4. When GPT-4o-mini used captions instead of images as input, there was a slight improvement across all metrics, but it did not show a significant difference compared to other open-source models. Notably, GPT-4o’s performance decreased in $exp_{\text{base}}$ but slightly improved in $exp_{\text{hint}}$ setting.
Model $exp_{\text{base}}$ $exp_{\text{hint}}$
Step(Mean.) Step(Std.) Step(Mean.) Step(Std.)
[…]
Table 11: Mean and standard deviation of gameplay steps for each model. The standard deviation was calculated using only successful escape trajectories. For models that failed to escape entirely, the standard deviation is marked with "-".
D.8 Reasoning Decompsition
The proportions of reasoning components and their composition for other models are shown in Table 12 and Figure 11.
Figure 11: Proportion for each reasoning component for other models.
Appendix E Detailed Explanation
E.1 Reasoning Component
We have classified the reasoning processes that models undertake while solving problems in VisEscape as follows: Observation, Guess, Hypothesis Formation, Planning, and Recall. The classification criteria are as follows:
  • Observation: Visually perceiving, describing, or analyzing the current scene.
  • Guess: Making a guess or assumption based on priors or knowledge, without exact evidence acquired in the environment.
  • Hypothesis Formation: Forming a hypothesis or idea to test in the environment, based on any evidence acquired in the environment.
  • Planning: Establishing or formulating action plans within the environment.
  • Recall: Retrieving information from history - remembering the state/location of objects, actions previously performed, and other relevant details.
We extract the pre-action reasoning and action performed by each model at every game step from the experimental logs, which were run with an applied memory management and a reasoning module. And they are annotated using GPT-4o achiam2023gpt. We set temperature as 0.0 for reproduction. To ensure the reliability of our machine-based evaluation method, we sampled 100 instances for each of the categories classified by GPT-4o, totaling 500 samples. These sentences were then manually annotated by humans following the same procedure. Figure 12 shows a heatmap between human and GPT-4o annotations, suggesting that while the model’s classifications largely align with human annotations, the model frequently misclassified the Guess type as Hypothesis Formation.
Figure 12: Heatmap of agreement between GPT-4o and human annotations for reasoning type classification. In 85.6% of the instances, the category annotations from GPT-4o and humans were identical.
Model Composition Ratio Length
[…]
Table 12: Reasoning composition for other models.
E.2 Memory Management Module
  • Structured spatial memory: This memory hierarchically stores information about room walls, the receptacles on those walls, the states of these receptacles, along with their visual information and characteristics. This serves as a directional aid for navigating toward specific receptacles or items.
  • Exploration memory: This memory stores information about which objects and items have been inspected so far and which have not. This encourages the agent to explore unobserved information rather than getting stuck on the information acquired thus far.
  • Salient action memory: This memory selectively stores actions that are worth remembering, such as interactions with objects, attempting to enter passwords, etc. This allows the agent to make action decisions by referencing actions it has performed in the past.
E.3 Visualization of Interactions in VisEscape
Figure 13 is a visualization of interactions defined in VisEscape. The size of each word is proportional to its frequency of appearance within the game.
Figure 13: Wordcloud of all types of actions defined in VisEscape. We excluded “open” and “close”, which are available on most receptacles.
E.4 Example of Visual Quizzes
Figure 14 shows two examples of visual quizzes in VisEscape.
Figure 14: Examples of visual quizzes that requires associative thinking in VisEscape. For all problems, we controlled the spatial positioning of the locks and clue images to ensure they remain distinctly separated. This design forces the agent not only to store each piece of information independently but also to identify relevant connections between them (such as “color” in the example above) and reason based on the established associations between these separate pieces of information.
Appendix F Related works on Escape Rooms
EscapeBench qian2024escapebench proposes a benchmark to evaluate creative reasoning in text-only environments, with a focus on tool use as a proxy for measuring creativity. IDEA he2024ideaenhancingrulelearning presents an agent that mimics the loop of human rule-learning by integrating multiple forms of reasoning to guide its actions, focusing on solving text-based puzzles. MM-Escape wang2025multimodallargelanguagemodels aims to extend the escape room paradigm by building an embodied environment for the evaluation of MLLM agents. Departing from embodied setups, our visually grounded benchmark blends tool use with puzzle-solving to create a more flexible yet challenging reasoning framework for evaluating MLLMs.
Appendix G Human Study
Room SR(%) GC(%) Steps Duration(s)
1 75.0 92.5 41.8 231.3
2 100.0 100.0 67.3 319.5
3 50.0 87.5 71.1 319.5
4 75.0 93.8 112.0 447.0
5 100.0 100.0 77.3 274.5
6 100.0 100.0 37.0 117.0
7 100.0 100.0 46.8 225.0
8 100.0 100.0 53.0 567.3
9 75.0 90.0 52.0 567.3
10 75.0 91.7 54.0 204.5
11 50.0 79.2 37.8 231.3
12 75.0 92.9 31.3 144.0
13 100.0 100.0 41.5 199.8
14 100.0 100.0 57.3 468.3
15 100.0 100.0 36.0 151.0
16 100.0 100.0 40.8 106.0
17 25.0 96.9 63.3 488.3
18 100.0 100.0 40.5 399.3
19 75.0 91.7 43.8 378.5
20 75.0 100.0 50.8 530.3
Average 82.5 95.8 52.8 318.5
Table 13: Human evaluation per-room success rates, goal completions, step counts, and time durations.
To compare the performance of AI models and human performance on escape room games, we conducted a human study. We recruited a total of 20 participants, including both individuals familiar and those unfamiliar with escape room games, and each was instructed to complete tasks in four different rooms. Figures 23 illustrate the Gradio interface used to provide the UI for playing VisEscape.
Table 13 presents the results of the human evaluation for 20 rooms, including the success rate, goal completion, number of steps, and time duration for experiments. Participants achieved high success rates in most rooms. We paid them based on the minimum hourly wage.
Variables:
direction, available_actions
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
Prompt:
{Initial Prompt}
<Current Observation>{direction} side of room - [IMAGE]</Current Observation>
Based on these information, choose next action to progress in the game. You can do one of the following actions: {available_actions}
Before you act, think first, and then act. Your thought should be in section [THINK], and your action should be in section [ACTION]. In [ACTION], respond ONLY with the chosen action, no other text.
[THINK]
[ACTION]
(a) Prompts for Reasoning module on initial step
Variables:
memory, salient_action_history, action_history, direction, current_scene_desc, previous_react, available_actions, hint_guideline_text
Puzzle text [OPTIONAL: if ispuzzle mode]:
<ANSWER> is an action to input the answer to open the lock you are facing. When you choose <ANSWER>, you should follow this format: <̈ANSWER>your answer</ANSWER>.̈
Prompt (Reason):
{Initial Prompt}
<Memory>{memory}</Memory>
<Action Memory>{salient_action_history}</Action Memory>
<Recent actions(from oldest to latest)>{action_history}</Recent actions>
<Current Observation>{direction} side of room - {current_scene_desc}</Current Observation>
<Your Thought and Action before this turn>{previous_react}</Your Thought and Action before this turn>
{available_actions}
{Puzzle text}
Before you act, think first, and then act. If there is a hint message, you should choose action to accomplish the guideline in hint message. {hint_guideline_text}
Your thought should be in section [THINK], and your action should be in section [ACTION]. In [ACTION], respond ONLY with the chosen action or <ANSWER>your answer</ANSWER>, no other text.
Prompt (Retry):
{Initial Prompt}
<Memory>{memory}</Memory>
<Your Previous Action> {before_action}
<Available Actions> {available_actions}
You just performed the <Your Previous Action>, but that action is not currently available in <Available Actions>. Referring to your memory, choose an action that is necessary to perform <Your Previous Action>.
{hint_guideline_text}
You should choose one of the following actions:{available_actions}
Please respond following below format without any other text: [ACTION]
(b) Prompts for Reasoning module and Retry
Table 14: Prompts for the Reasoning module-before action.
Variables:
action_history
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
last_10_history (for log in action_history) :
[…]
Action: [{log[’action’]}]-{log[’analysis’]}
spatial_json_format:
{"direction 1" : {
   "objects":["object1", "object2", ...]
inspected_objects_json_format:
[{"object 1" : {
   "state":"",
   "characteristics":"",
   "additional info":""
Prompt:
{Initial Prompt}
<Last 10 logs(from oldest to latest)> {last_10_history} </Last 10 logs>
Here is the definition of information:
<Last 10 logs>: Sequence of observation, action-effect, next observation, next action-effect for each turn.
Here is the task:
Construct your memory about the room, based on the last 10 logs. Follow below guidelines:
1. Identify which objects exist on each directional wall, and add the information to "[SPATIAL MEMORY]" section.
2. If you have inspected an object via "inspect [object]", you should add the information to "[INSPECTED OBJECTS]" section.
3. The objects that you have not inspected via "inspect [object]" should be added to "[UNINSPECTED OBJECTS]" section.
4. Add any information from observations which is not included in [SPATIAL MEMORY] and [INSPECTED OBJECTS] that you think is necessary for solving other problems to the "[Additional Memory]" section.
Please respond following below format without any other text:
[SPATIAL MEMORY] {spatial_json_format}
[INSPECTED OBJECTS] {inspected_objects_json_format}
[UNINSPECTED OBJECTS] []
[ADDITIONAL MEMORY] [1. additional memory1, 2. additional memory2, ...]
Table 15: Prompts for memory management module - for first run (memory construction).
Variables:
spatial_memory, action_history
Initial prompt:
You are an AI agent playing a room escape game. The room is surrounded by 4 walls, and you can explore other walls by "turn_to_[direction]". Each wall has objects that you can interact with, and you can inspect the object by "inspect [object]". Please read the given information and task description below carefully and respond accordingly.
last_10_history (for log in action_history) :
[…]
Action: [{log[’action’]}]-{log[’analysis’]}
spatial_json_format:
{"direction 1" : {
   "objects":["object1", "object2", ...]
inspected_objects_json_format:
[{"object 1" : {
   "state":"",
   "characteristics":"",
   "additional info":""
Prompt:
{Initial Prompt}
<Current Memory> {spatial_memory} </Current Memory>
<Last 10 logs(from oldest to latest)> {last_10_history} </Last 10 logs>
Here is the definition of each information:
<Last 10 logs>: Sequence of observation, action-effect, next observation, next action-effect for each turn.
Here is the task:
Update your memory about the room, based on the last 10 logs. Follow below guidelines:
1. If you newly inspected an object via "inspect [object]" among objects in "[UNINSPECTED OBJECTS]" section,
  you should add the information to "[INSPECTED OBJECTS]" section.
2. Based on the information that you can obtain from last 10 logs, update <Current Memory>
  if there exists any information that you can obtain from <Last 10 logs> but not in <Current Memory>.
3. Add any information not included in spatial memory and inspected objects that you think is necessary
  for solving other problems to the "[ADDITIONAL MEMORY]" section.
Please respond following below format without any other text:
[SPATIAL MEMORY] {spatial_json_format}
[INSPECTED OBJECTS] {inspected_objects_json_format}
[UNINSPECTED OBJECTS] []
[ADDITIONAL MEMORY] [1. additional memory1, 2. additional memory2, ...]
Table 16: Prompts for Memory management module after memory construction
Variables:
previous_scene, previous_action, current_scene
Prompt:
{Initial Prompt}
<Previous Observation> : {previous_scene}
<Previous Action> : {previous_action}
<Current Observation> : {current_scene}
Here is the definition of each information:
<Previous Observation> is a description of a scene before your action, and <Current Observation> is a description of a scene after your action.
Here is the task:
Analyze the effect of your action by comparing <Previous Observation> and <Current Observation>. The analysis should be concise and definitive, not descriptive.
Keep it under 10 words.
[ANALYSIS]
Table 17: Prompts for Reasoning module-after action.
Variables:
item_name
Prompt:
This image is a close-up view of an item ’{item_name}’.
Describe this image. Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(a) Prompts for getting caption from image of item-view
Variables:
object_type, items_str
This image is a close-up view of an object ’{object_type}’.
In {object_type}, the following objects are present: {items_str}
Describe this image. The names of visible objects should be expressed using the given object names above, enclosed in "".
Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not include any analysis of the scene or the room, just describe the image.
3. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(b) Prompts for getting caption from image of object-view
Variables:
objects
This image is a wall view of a room, with following objects: {objects}.
Describe this image. The names of visible objects should be expressed using the given object names above, enclosed in "".
Your description should fulfill the following rules:
1. Description should include every visual information, but concise and clear.
2. Do not include any analysis of the scene or the room, just describe the image.
3. Do not start description with phrases like ’The image depicts’, ’The image shows’, etc.
(c) Prompts for getting caption from image of wall-view
Table 18: Prompts for captioning observations.
Prompt:
You are a game designer proposing ideas for an ’Escape Room Game’. I want to create an ’Escape Room Game’ in a room structure enclosed by four walls. Inside the room, there are various objects, and it’s an ’Escape Room Game’ where actions initiated from a specific object trigger other objects, allowing puzzles to be solved one by one. There are two types of objects: 1. Receptacle:
- Definition) Objects that can contain or hold other objects, such as tables, drawers, or wardrobes.
- Each Receptacle is visually distinct and has unique States. For example, a door without a lock has two States: ’open’ and ’closed’, but a door with a lock has three States: ’locked’, ’open’, and ’closed’. Also, a drawer with two compartments can have four States depending on the open/closed status of each compartment: ’opened_opened’, ’opened_closed’, ’closed_opened’, ’closed_closed’.
- Each Receptacle must be assigned to one of the four walls of the room.
- Receptacles can contain Items. However, considering realistic constraints, each Item must define the "Interactable State(s)" of the Receptacle where it can be interacted with (i.e., visible). For example, if you have a key and a wardrobe with States defined as [locked, closed, open], the key’s interactable state for that wardrobe is [open]. Interactable States can be multiple. For instance, for a drawer with two compartments, if a key exists in the second compartment, the key’s interactable states are [closed_open, open_open].
2. Item:
- Definition) Objects that can be triggers for interaction with specific Receptacles and other Items, such as keys, buttons, photos, etc. Items must be contained within a Receptacle. There are two types of Items:
- 2-a. Applicable Item: Items that can be picked up, stored in an inventory, and applied to other objects, such as keys or ID cards.
- 2-b. Quiz Item: Items that contain a puzzle, which, when correctly solved or triggered by a specific action, activates a change in another specific object within the game.
Furthermore, in addition to objects, the game requires internal logic for escaping the room. For example, a wardrobe is initially locked, but finding and applying a key from elsewhere unlocks it. A drawer is initially locked, but solving a quiz elsewhere opens it. The game logic can be very creative and diverse, but it must be defined according to the following constraints: 1. It must be limited to logic that changes the state of a Receptacle.
2. Actions related to game logic must be associated with an Item. For example, changing a drawer from a ’locked’ State to a ’closed’ State via an ’apply Key’ action is game logic. However, changing a drawer from an ’opened’ State to a ’closed’ State via a ’close’ action is not game logic.
If these two constraints are met, the in-game logic can be expressed atomically in the following format:
(receptacle_name.state_name_before_action, action_name(item_name),
receptacle_name.state_name_after_action, checkTrigger(interaction_name)) In summary, to design one escape room game, the following game components are needed:
- Approximately 5 Receptacles, all possible states defined for each object, and which wall they belong to.
- Approximately 5 Items, and which Receptacle they belong to.
- Approximately 5 game logic interactions, expressed in the atomic format described above.
Please design a game logic with following receptacles and items:
[Receptacle] - 2TierDrawer, Cupboard, Safe, Wardrobe, Desk, Door
[Receptacle] - Key1, key2, ID card, button, puzzle
Design an escape room game by creating interactions between objects and puzzles. Storytelling is not required, and the game does not need to be particularly creative or innovative.
Table 19: Prompts used for game logic design.
Variables:
image1, image2, [Optional]tried_answers
{History} (if tried_answers is given):
You have tried the following answers: {tried_answers}, but they are all wrong.
Prompt:
{image1} is a hint image for the password of the lock described in {image2}.
You should guess the password of the lock solely based on the hint image.
{History}
Respond ONLY with the your answer, no other text.
(a) Prompts for multi-image VQA without reasoning.
Variables:
image1, image2, [Optional]tried_answers
{History} (if tried_answers is given):
You have tried the following answers: {tried_answers}, but they are all wrong.
Prompt:
{image1} is a hint image for the password of the lock described in {image2}.
You should guess the password of the lock solely based on the hint image.
{History}
Before you guess, think first, and then guess the password. Your thought should be in section [Think], and your answer should be in section [Answer]. In [Answer], respond ONLY with the your answer, no other text.
[Think]
[Answer]
(b) Prompts for multi-image VQA with reasoning.
Table 20: Prompts for multi-image VQA.
Variables:
think, action
Prompt:
You are tasked with analyzing the logs of an AI agent in an escape room game.
I’ll give you a think-action pair that the AI agent has performed.
Your task is as follows:
1. Divide the agent’s reasoning chain into distinct units that fall into the categories explicitly given below:
- Observation: visually perceiving, describing, or analyzing the current scene
- Guess: making a guess or assumption based on priors or knowledge, without exact evidence acquired in the environment
- Hypothesis Formation: forming a hypothesis or idea to test in the environment, based on any evidence acquired in the environment
- Planning: Establishing or formulating action plans within the environment
- Recall: Retrieving information from history - remembering the state/location of objects, actions previously performed, and other relevant details
A unit could be:
- A single sentence
- Multiple sentences together
- Multiple units within a single sentence
2. If a unit does not fit into the above categories, create a new category called [None-<new category name>] and place it there.
3. List all units in the same order as they appear in the original text. Do not reorder or rearrange the units.
Input:
[Think]{think}
[Action]{action}
Response Format:
[Category of Unit 1]: Sentence or phrase of given [Think]
[Category of Unit 2]: Sentence or phrase of given [Think]
[Category of Unit 3]: Sentence or phrase of given [Think]
Response:
Table 21: Prompts for annotating components of reasoning.
Figure 22: Example of rooms rendered using Chaos Enscape.
Figure 23: User Interface generated by Gradio for conducting experiments for human trajectory.